Circular No. 8640

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005lb–2005lq CBET 315 contains information for sixteen newly discovered super-

novae found on multiple g, r, and i images taken with the SDSS 2.5-m telescope and reported by J. Frieman on behalf of the Sloan Digital Sky Survey II collaboration; all of the new objects are type-Ia supernovae except for 2005lb and 2005lc (type-II events) and 2005lm (probable type-II event). Fourteen of the supernovae were discovered at mag g > 20; the other two are tabulated below:

SN  2005 UT  α_{2000}  δ_{2000}  Mag. z
2005lb Sept. 10 22^{h}54^{m}50.^{s}05 − 0^{o}15′09′′.0 18.0 0.03
2005lc Sept. 10 3 02 11.18 − 1 09 59.4 18.8 0.01

V1663 AQUILAE R. C. Puetter, University of California at San Diego; R. J. Rudy, D.

K. Lynch, S. Mazuk, and C. C. Venturini, The Aerospace Corporation; R. B. Perry, Langley Research Center, NASA; and B. Walp, Lick Observatory, report 0.47- to 2.5-μm spectroscopy of V1663 Aql (cf. IAUC 8540, 8544) with the Lick 3.0-m telescope (+ VNIRIS) at Nov. 14.160 UT. The optical spectrum was largely nebular, showing [O III], [N II], Hα, and weak He I 587.6-nm, He I 706.5-nm, and [O I]. Weak neutral lines of O I and C I were present. The spectrum showed coronal-line emission of [S VIII], [S IX], strong [Si VI] and [Si VII], weak [Ca VIII], and the unidentified novae lines. The emission lines were flat-topped with minor structure, and their FWHM and FWZI were 2000 km/s and 2600 km/s, respectively. Based on the O I lines, which are about equally produced by Lyβ and continuum fluorescence, the reddening is possibly as large as 2 magnitudes in E(B−V).

COMETS C/2005 S12 AND C/2005 S13 (SOHO) Two additional Kreutz sungrazing comets have been found on SOHO

website images (cf. IAUC 8638), their first available positions given below. C/2005 S12 was stellar in C3 images, reaching mag 7.0 at 11.1R_{⊙} on Sept. 26.431. S/2005 S13 was tiny, stellar, and too faint for photometry in C3 images, and both objects were faint and diffuse in C2 images.

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 S12 Sept. 26.154 11^{h}53.^{m}8 − 2^{o}05′ C3/2 HS 2005-X11
C/2005 S13  29.821 12 05.8 − 3 41 C3/2 BZ 2005-X11
2005 December 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green